OpenStudy (anonymous):

-4 1/5 -(-2 3/5)= explain....

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Nnesha

TheSmartOne (thesmartone):

when you subtract a negative number it becomes addition..

OpenStudy (anonymous):

well now i have a new problem cause it took forever for someone to come on here , no one would explain it :d

TheSmartOne (thesmartone):

\[-4\frac{ 1 }{ 5 }- (-2\frac{ 3 }{ 5 })=-4\frac{ 1 }{ 5}+2\frac{ 3 }{ 5}\]

OpenStudy (anonymous):

hold on i have a new one

OpenStudy (anonymous):

3 1/5 x (-1 2/5)

OpenStudy (anonymous):

would it be -112x/25??

OpenStudy (anonymous):

??

TheSmartOne (thesmartone):

\[3\frac{ 1 }{ 5 }\times -1\frac{ 2 }{5}=\frac{ 16 }{ 5 }\times \frac{ -7 }{5 }=\frac{ -112 }{ 25 }\]

OpenStudy (anonymous):

oh haha without the x i see and wait

OpenStudy (anonymous):

it said it was wrong O.o

OpenStudy (anonymous):

OMG every answer is wrong?!?!?!

OpenStudy (anonymous):

-2/3 x (-1/6)

TheSmartOne (thesmartone):

\[\frac{ -2 }{ 3 }\times \frac{ -1 }{ 6 }=\frac{ 2 }{ 18 }=\frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

1 2/5 x 4/5?

TheSmartOne (thesmartone):

You try it -_-

OpenStudy (anonymous):

then can you check my answer if it is right or not??

TheSmartOne (thesmartone):

sure

OpenStudy (anonymous):

28/25?

TheSmartOne (thesmartone):

Correct. ^_^

OpenStudy (anonymous):

O.o NO

OpenStudy (anonymous):

it said it was wrong!

OpenStudy (anonymous):

OMG im about to break this computer